A 500 m3 living room is to be heated by an electric heater. It is expected that the heater be able to increase the air temperature in the room from 5C to 25C in 1 hour. Assuming no heat losses from the sealed room and an atmospheric pressure of 100kPa, (a) determine the required power in kW of the resistance heater? (b) The total cost of the electrical energy consumed during this 1-hour period. Take the unit cost of electricity to be TL0.80/kWh ?
1
Expert's answer
2020-11-16T07:47:44-0500

a)


"Pt=Q=mc(T_f-T_i)=\rho Vc(T_f-T_i)\\3600P=(1.225) (500)(0.718)(25-5)\\P=2.44\ kW"

b)


"S=2.44(0.80)=\$\ 1.95"